determine the proper sign and acute angle (0
Answer by Alan3354(69443)

determine the proper sign and acute angle (0 23pi/6 = 24pi/6 - pi/6 = -pi/6
sin(-pi/6) = -1/2
--------------
It's not in Quadrant 1, I'm not sure what you mean by the rest.